OpenStudy (legomyego180):

Find a formula for the sequence:

OpenStudy (legomyego180):

\[\left\{ -1, \frac{ 2 }{ 3 }, -\frac{ 1 }{ 3 }, \frac{ 4 }{ 27 },- \frac{ 5 }{ 81 } \right\}\]

OpenStudy (legomyego180):

I understand the alternating negatives, but the third term is throwing me off as far as any relationship between the numerators

OpenStudy (legomyego180):

Would it help to find the LCD here?

OpenStudy (legomyego180):

@jim_thompson5910 @zepdrix @AloneS @pooja195

OpenStudy (legomyego180):

I rewrote this as: \[\left\{- \frac{ 81 }{ 81 }, \frac{ 54 }{ 81 }, -\frac{ 27 }{ 81 }, \frac{ 12 }{ 81 }, -\frac{ 5 }{ 81 } \right\}\] Didnt really help though

jimthompson5910 (jim_thompson5910):

\[\Large \left\{ -1, \frac{ 2 }{ 3 }, -\frac{ 1 }{ 3 }, \frac{ 4 }{ 27 },- \frac{ 5 }{ 81 } \right\}\] is equivalent to \[\Large \left\{ -\frac{1}{1}, \frac{ 2 }{ 3 }, -\frac{ 3 }{ 9 }, \frac{ 4 }{ 27 },- \frac{ 5 }{ 81 } \right\}\] I changed the -1 to -1/1 I changed the -1/3 to -3/9

OpenStudy (legomyego180):

Oh that looks a lot better

jimthompson5910 (jim_thompson5910):

the numerators (1,2,3,4,5) are incrementing by 1 the denominators are powers of 3 1 = 3^0 3 = 3^1 9 = 3^2 27 = 3^3 81 = 3^4 also, the terms are alternating in sign, so you'll need a (-1)^n or (-1)^(n+1) term in there

OpenStudy (legomyego180):

so (n-1)^3 for the denominator?

OpenStudy (legomyego180):

no, thats not right. The denominator is confusing me, Im not seeing how to relate it to n.

jimthompson5910 (jim_thompson5910):

no, not cubed, powers of 3

jimthompson5910 (jim_thompson5910):

x^3 is different from 3^x

jimthompson5910 (jim_thompson5910):

the denominator would be 3^(n-1) if you start n at n = 1

OpenStudy (legomyego180):

\[a _{n}=\frac{ n }{ 3^{(n-1)} }(-1^{n})\]

OpenStudy (legomyego180):

I think this is the correct answer

jimthompson5910 (jim_thompson5910):

it's very close, but not quite there

jimthompson5910 (jim_thompson5910):

you need to write (-1^n) as (-1)^n

OpenStudy (legomyego180):

n needs to be outside the parentheses?

OpenStudy (legomyego180):

sweet ok, gotcha. Thanks for your help Jim

jimthompson5910 (jim_thompson5910):

consider when n = 2 (-1^n) = (-1^2) = (-1*1^2) = -1 while (-1)^n = (-1)^2 = (-1)*(-1) = 1

jimthompson5910 (jim_thompson5910):

no problem